OpenStudy (anonymous):

Simplify 24over36 x 9over27 24/36x9/27 Please help?

OpenStudy (zzr0ck3r):

\[\frac{24}{36}*\frac{9}{27}*\frac{24}{36}*\frac{9}{27}\]?

OpenStudy (zzr0ck3r):

12 goes into 24 2 times and 12 goes into 36 3 times so \[\frac{24}{36}=\frac{2}{3}\]

OpenStudy (zzr0ck3r):

9 goes into 9 1 time and 9 goes into 27 3 times so \[\frac{9}{27}=\frac{1}{3}\]

OpenStudy (anonymous):

24/36 = 2/3 9/27 = 1/3 now you have 2/3 x 1/3

OpenStudy (zzr0ck3r):

so \[\frac{24}{36}*\frac{9}{27}*\frac{24}{36}*\frac{9}{27}=\frac{2}{3}*\frac{1}{3}*\frac{2}{3}*\frac{1}{3}=\frac{2*1*2*1}{3*3*3*3}=\frac{4}{81}\]

OpenStudy (anonymous):

24/36 * 9/27 = 2/3 * 1/3 = 2/9 or if you don't want to reduce at the beginning you can just multiply and reduce at the end 24/36 * 9/27 = 216/972 = 108/486 = 54/243 = 2/9